OpenStudy (abbycross167):

Can someone please check my work on this math question? please? I'll fan and give a medal!!

OpenStudy (abbycross167):

18. A tennis ball is launched with an initial velocity of 24.5 m/s from the edge of a cliff that is 117.6 meters above the ground. Which quadratic equation could be used to correctly determine when the ball will hit the ground: 4.9t2 + 24.5t + 117.6 = 0 -4.9t2 - 24.5t + 117.6 = 0 -4.9t2 + 24.5t - 117.6 = 0 4.9t2 + 24.5t - 117.6 = 0 -4.9t2 + 24.5t + 117.6 = 0

OpenStudy (abbycross167):

my answer is -4.9t2 + 24.5t + 117.6 = 0

OpenStudy (zzr0ck3r):

why?

OpenStudy (zzr0ck3r):

is this for a calc class?

OpenStudy (whpalmer4):

the equation of vertical position for something launched from initial height \(h_0\) with vertical velocity \(v_0\) and acceleration due to gravity \(g\) will be \[h(t) = -\frac{1}{2}gt^2 + v_0 t + h_0\] Your answer is correct, as \(g\) is usually taken as \(\approx 9.8 \text{ m/s}^2\) and a positive value for \(v_0\) means the ball is initially traveling upwards.

OpenStudy (zzr0ck3r):

From a math standpoint, we know the derivative give velocity. So the derivative needs to be something that we plug in 0 and get out 24.5 obviously the constant term is 117.6 do yes :)

OpenStudy (abbycross167):

thank you very much!!